# Canonical Electrodynamics in Ashtekar-Barbero variables

### v1

@@ 5 Discussion and conclusions
In this work we have provided the full analysis of the canonical formulation of the system of gravity, fermions, and photons, and their interaction in a background-independent way. In particular, we have considered the interaction term of the QED, which to our knowledge, has not been considered in previous studies.
We have analyzed the constraint structure of the system in the Ashtekar-Barbero variables, with either Weyl or Dirac fermions, and have shown how the connection variables are modified due to the presence of fermions. Furthermore, the expression for the Hamiltonian, spatial diffeomorphism and Guass constraints are given in a way that the contribution of each of the sectors, i.e., gravity, fermions, and photons, are clear. We have also derived the EOM of the matter sector, which reduce to the standard EOM of QFT on fixed curved or flat spacetime in the appropriate limits. The behavior of the system under parity transformation is studied in detail. Finally, we presented the modifications to the constraints arising from the introduction of non-minimal coupling between gravity and fermions.
This analysis and the results pave the way for further canonical analysis of realistic matter in the universe involved in electromagnetic interactions in strong and dynamical (non-fixed) gravity regimes. Most importantly, this work opens the door for further detailed studies of an interactive system (QED in a dynamical background) using loop quantum gravity, polymer quantization, generalized uncertainty principle, or other theory that uses canonical quantization techniques. A rather immediate application of this analysis is in symmetry-reduced systems such as cosmological models involving electromagnetic matter, or collapse scenarios in spherical symmetric or rotating black holes, which up to now, have been done almost exclusively only with scalar fields.
@@ Acknowledgements.
Appendix A Complementary information on fermions
Appendix A serves to clarify the notation adopted for the fermionic field, the gamma matrices, and their associated identities. Given the ambiguity that exists in the literature, this appendix establishes the conventions used throughout the paper to ensure consistency.
A.1 Notes on helicity and chirality
In this subsection, we introduce some key concepts related to helicity and chirality Pal:2011pb ; Schwartz:2014sze , which will help clarify the discussions that follow.
A general definition of “handedness” applicable to any particle involves the relative orientation between its momentum and angular momentum. For a particle with three-momentum $\mathbf{p}$, helicity is defined as
$h_{\mathbf{p}}\coloneqq\frac{2\mathbf{J}\cdot\mathbf{p}}{|\mathbf{p}|}\,,$ (159)
where $\mathbf{J}$ is the total angular momentum of the particle. Since the orbital angular momentum is always perpendicular to the direction of motion, it does not contribute to helicity. Therefore, helicity effectively measures twice the projection of the spin along the direction of momentum. Defining the spin matrices as $\frac{1}{2}\Sigma$, the helicity operator can be rewritten as
$h_{\mathbf{p}}\coloneqq\frac{\Sigma\cdot\mathbf{p}}{|\mathbf{p}|}\,.$ (160)
It follows that the eigenvalues of $h_{\mathbf{p}}$ are ±$\pm 1$. An eigenstate with eigenvalue $-1$ is referred to as “left-handed”, while one with eigenvalue $+1$ is called “right-handed”.
A key property of helicity is that it commutes with the Dirac Hamiltonian, which means it is conserved in coordinate time. Moreover, helicity is invariant under rotations, as evident from the rotational invariance of the dot product in its definition. In other words, different observers related by a spatial rotation will agree on the helicity of a given particle. However, helicity is not invariant under Lorentz boosts. As a result, a massive fermion [^3] The change in helicity can only occur in frames moving faster than the particle. Since such a frame cannot exist for a massless particle, helicity remains invariant in that case. cannot be strictly left- or right-helical in an absolute sense; the helicity depends on the observer’s frame of reference.
With regard to the definition of chirality, it is associated with the matrix $\gamma_{5}$ (see Appendix A.2 for its properties). This matrix allows us to construct projection operators that act on fermion fields as follows
$\mathds{L}=$ $\frac{1}{2}\left(1-\gamma_{5}\right),$ $\mathds{R}=$ $\frac{1}{2}\left(1+\gamma_{5}\right)\,.$ (161)
These projection operators are commonly referred to as “left-chiral” and “right-chiral”, respectively. Given any Dirac spinor $\Psi(x)$ that satisfies the Dirac equation, we can decompose it into its left- and right-chiral components as
$\Psi=\Psi_{L}+\Psi_{R},$ (162)
@@ A.1 Notes on helicity and chirality
$\Psi_{L}=$ $\mathds{L}\Psi,$ $\Psi_{R}=$ $\mathds{R}\Psi\,.$ (163)
An important feature of chirality is that a left-chiral solution to the Dirac equation remains left-chiral under Lorentz transformations. This is ensured by the fact that $\gamma_{5}$ anticommutes with all gamma matrices, allowing the chiral projections to be defined in a Lorentz covariant manner. However, chirality is not conserved for a massive free particle, as $\gamma_{5}$ does not commute with the mass term in the Dirac Hamiltonian. Since the mass term involves only one gamma matrix, it anticommutes with $\gamma_{5}$, leading to a mixing of chiral components. In contrast, there is no such issue with the kinetic (derivative) term, which preserves chirality.
In this sense, chirality and helicity exhibit complementary behaviors: helicity is conserved for a free particle but is not Lorentz invariant, whereas chirality is Lorentz invariant but not conserved when the particle has mass.
In the massless case, applying the Dirac equation to a spinor eigenstate of 3-momentum $\mathbf{p}$, we find
$\gamma_{5}w_{\mathbf{p}}=\frac{\Sigma\cdot\mathbf{p}}{|\mathbf{p}|}w_{\mathbf{p}}\,,$ (164)
where $w_{\mathbf{p}}$ denotes a general spinor. This relation shows that chirality and helicity coincide for massless spinors: the action of $\gamma_{5}$ on the spinor yields its helicity. In this limit, chirality becomes a physical observable and Lorentz-invariant quantity.
When mass is introduced, the equations of motion cause mixing between left- and right-chiral components. However, since momentum and spin remain good quantum numbers in the free theory —even for massive particles— helicity is still conserved. As a result, helicity remains a useful and meaningful concept in the massive case. The crucial distinction is that, in the presence of mass, helicity eigenstates no longer coincide with chirality eigenstates.
A.2 Weyl fermions in flat spacetime
In order to maintain consistency with the majority of the literature regarding fermionic matter Schwartz:2014sze ; Weinberg:1995s ; Peskin:1995ev , we will switch to the standard QFT signature convention, (+)(+\,-\,-\,-)( + - - - ), for Section A.2 only.
In general, elements of the vector space on which the spin-$\frac{1}{2}$ representations act are known as spinors. When we consider massless fermions, we can talk of positive or negative helicity fermion or of left or right chiral fermion  [^4] In the massless case, helicity are chiraly are related to each other (see Appendix A.1).. It is well known that the proper Lorentz algebra $\mathfrak{so}(1,3)$ is isomorphic to the Lie algebra of $SU(2)\times SU(2)$, meaning that any representation of the Lorentz algebra can be characterized by how it transforms under each of the two $\mathfrak{su}(2)$ factors. In this framework, a left-chiral fermion transforms as a doublet under one $\mathfrak{su}(2)$ and as a singlet under the other, denoted by the representation $(\frac{1}{2},0)$. Conversely, a right-chiral fermion transforms as $(0,\frac{1}{2})$. These are known as Weyl fermions Pal:2011pb ; Schwartz:2014sze
$\psi_{R}=$ $\begin{pmatrix}\psi_{1}(x)\\ \psi_{2}(x)\end{pmatrix},$ for the (0,12) $\text{for the $\left(0,\frac{1}{2}\right)$ representation}\,,$ (165)
$\psi_{L}=$ $\begin{pmatrix}\psi_{1}(x)\\ \psi_{2}(x)\end{pmatrix},$ for the (12,0) $\text{for the $\left(\frac{1}{2},0\right)$ representation}\,.$ (166)
Under rotational angles $\theta_{i}$ and boost angles $\omega_{i}$, they transform as
$\psi_{L}\rightarrow\Lambda_{L}\psi_{L}=$ $e^{i\frac{\bar{\sigma}}{2}\left(\bar{\theta}+i\bar{\omega}\right)}\psi_{L},$ $\psi_{R}\rightarrow\Lambda_{R}\psi_{R}=$ $e^{i\frac{\bar{\sigma}}{2}\left(\bar{\theta}-i\bar{\omega}\right)}\psi_{R}\,.$ (167)
Since $\Lambda_{L}$ and $\Lambda_{R}$ are complex, so are $\psi_{L}$ and $\psi_{R}$ Similarly,
$\psi_{L}^{\dagger}\rightarrow$ $\psi_{L}^{\dagger}e^{-i\frac{\bar{\sigma}}{2}\left(\bar{\theta}-i\bar{\omega}\right)},$ $\psi_{R}^{\dagger}\rightarrow$ $\psi_{R}^{\dagger}e^{-i\frac{\bar{\sigma}}{2}\left(\bar{\theta}+i\bar{\omega}\right)}\,.$ (168)
By definition, Weyl fermions are massless, so we are only concerned with the kinetic term of their Lagrangian. To derive this term, consider the vector quantities $v^{i}=\psi_{R}^{\dagger}\sigma^{i}\psi_{R}$ and $\tilde{v}^{i}=\psi_{L}^{\dagger}\sigma^{i}\psi_{L}$, where $\sigma^{i}$ denotes the $i$th Pauli matrix. Using the well-known identities of the Pauli matrices,
$\left[\sigma^{i},\sigma^{j}\right]=$ $2i\epsilon^{ij}\,_{k}\sigma^{k}\,,$$\left\{\sigma^{i},\sigma^{j}\right\}=$ $2\delta^{ij}\mathbb{1}\,,$ (169)
we can compute the infinitesimal transformations of $v^{i}$ and $\tilde{v}^{i}$ under Lorentz transformations. This yields
$\delta v^{i}=$ $-\epsilon^{ij}\,_{k}\theta_{j}\psi_{R}^{\dagger}\sigma^{k}\psi_{R}+\omega^{i}\psi_{R}^{\dagger}\psi_{R}\,,$ $\delta\tilde{v}^{i}=$ $-\epsilon^{ij}\,_{k}\theta_{j}\psi_{L}^{\dagger}\sigma^{k}\psi_{L}-\omega^{i}\psi_{L}^{\dagger}\psi_{L}\,.$ (170)
Thus, both $v^{i}$ and $\tilde{v}^{i}$ transform as the spatial parts of a Lorentz 4-vector $\delta V^{\mu}=\Lambda_{\nu}^{\mu}V^{\nu}$ with $\Lambda_{i}^{0}=\omega^{i}$ and $\Lambda_{j}^{i}=\epsilon^{i}\,_{jk}\theta^{k}$. We can define
$v^{\mu}=$ $(\psi_{R}^{\dagger}\psi_{R},\psi_{R}^{\dagger}\sigma^{i}\psi_{R})\,,$ $\tilde{v}^{\mu}=$ $(\psi_{L}^{\dagger}\psi_{L},-\psi_{L}^{\dagger}\sigma^{i}\psi_{L})\,,$ (171)
so that the combinations $\psi_{R,L}^{\dagger}\partial_{t}\psi_{R,L}$ transform as Lorentz invariants.
Defining the matrices
$\sigma^{\mu}=$ $(\mathbb{1},\sigma^{i})\,,$ $\tilde{\sigma}^{\mu}=$ $(\mathbb{1},-\sigma^{i})\,,$ (172)
we can write the previous vectors in the following form
$v^{\mu}=$ $\psi_{R}^{\dagger}\sigma^{\mu}\psi_{R}\,,$ $\tilde{v}^{\mu}=$ $\psi_{L}^{\dagger}\tilde{\sigma}^{\mu}\psi_{L}\,,$ (173)
and the kinetic terms of the Lagrangian become
$\mathcal{L}_{R}=$ $i\psi_{R}^{\dagger}\sigma^{\mu}\partial_{\mu}\psi_{R}\,,$ $\mathcal{L}_{L}=$ $i\psi_{L}^{\dagger}\tilde{\sigma}^{\mu}\partial_{\mu}\psi_{L}\,.$ (174)
Hence, the Dirac Lagrangian for Weyl fermions can be expressed as
$\mathcal{L}_{\text{Dirac}}=i\psi_{R}^{\dagger}\sigma^{\mu}\partial_{\mu}\psi_{R}+i\psi_{L}^{\dagger}\tilde{\sigma}^{\mu}\partial_{\mu}\psi_{L}\,.$ (175)
It is now possible to introduce a notation that can be used for arbitrary fermionic field combining the two spinors into a four-component object
$\Psi=$ $\begin{pmatrix}\psi_{L}\\ \psi_{R}\end{pmatrix},$ $\Psi^{\dagger}=$ $\begin{pmatrix}\psi_{L}^{\dagger}&\psi_{R}^{\dagger}\end{pmatrix}.$ (176)
In this basis, known as the Weyl (or chiral) representation, the gamma matrices take the form
$\gamma^{\mu}=\begin{pmatrix}0&\sigma^{\mu}\\ \tilde{\sigma}^{\mu}&0\end{pmatrix}\,,$ (177)
with components
$\gamma^{0}=$ $\begin{pmatrix}0&\mathbb{1}\\ \mathbb{1}&0\end{pmatrix}\,,$ $\gamma^{i}=$ $\begin{pmatrix}0&\sigma^{i}\\ -\sigma^{i}&0\end{pmatrix}\,.$ (178)
These matrices satisfy the Clifford algebra $\{\gamma^{\mu},\gamma^{\nu}\}=2\eta^{\mu\nu}$.
Moreover, we need to identify a suitable way to combine fermionic fields in order to construct a Lorentz scalar. This is achieved by introducing the Dirac adjoint (or conjugate) of the fermionic field, defined using the zeroth gamma matrix
$\overline{\Psi}=\Psi^{\dagger}\gamma^{0}=\begin{pmatrix}\psi_{L}^{\dagger}&\psi_{R}^{\dagger}\end{pmatrix}\begin{pmatrix}0&\mathbb{1}\\ \mathbb{1}&0\end{pmatrix}=\begin{pmatrix}\psi_{R}^{\dagger}&\psi_{L}^{\dagger}\end{pmatrix}\,.$ (179)
This definition ensures that the bilinear form $\overline{\Psi}\Psi$ transforms as a Lorentz scalar, and that the kinetic term $\overline{\Psi}\gamma^{\mu}\partial_{\mu}\Psi$ transforms as a Lorentz scalar density, preserving the overall covariance of the Lagrangian. Thus, we arrive at the conventional form of the Dirac Lagrangian for a Weyl fermionic field
$\mathcal{L}_{\text{Dirac}}=i\overline{\Psi}\gamma^{\mu}\partial_{\mu}\Psi\,.$ (180)
A.3 Weyl fermions in curved spacetime
Consider the fermionic field in a curved spacetime. In this context, it is necessary to introduce a suitable covariant derivative that accounts for the coupling between fermions and gravity. This is achieved by incorporating the generalized connection 1-form $\omega_{a}^{IJ}$. Hence, The covariant derivatives acting on the spinor field $\Psi$ and its conjugate field $\overline{\Psi}$ are defined respectively as
$\mathfrak{D}_{\mu}\Psi=$ $\partial_{\mu}\Psi+\frac{1}{2}\omega_{\mu}^{IJ}\sigma_{IJ}\Psi\,,$ (181)
$\overline{\mathfrak{D}_{\mu}\Psi}$ $=\partial_{\mu}\overline{\Psi}+\frac{1}{2}\omega_{\mu}^{IJ}\Psi^{\dagger}\sigma_{IJ}^{\dagger}\gamma^{0}\,,$ (182)
where $\sigma^{\mu\nu}\coloneqq\frac{1}{4}[\gamma^{\mu},\gamma^{\nu}]$ are the generators of the Lorentz group in the spinor representation. Using this formalism, the action for a Weyl fermion in curved spacetime is given by
However, in General Relativity we adopt the mostly-plus metric signature (,+,+,+)(-,+,+,+)( - , + , + , + ), which requires a modification of some algebraic relations involving the gamma matrices. While the Pauli matrices and their commutation and anticommutation relations remain unchanged, the Clifford algebra for the gamma matrices must be adapted accordingly. We choose to retain the explicit form of the gamma matrices as given in Eq.(178). However, due to the change in metric signature, the anticommutation relation now reads:
$\left\{\gamma^{I},\gamma^{J}\right\}=-2\eta^{IJ}\,,$ (184)
where $\eta^{IJ}$ is the Minkoswki metric. This relation leads to the following properties for the gamma matrices:
$\left(\gamma^{0}\right)^{\dagger}=\gamma^{0}$ $\implies\left(\gamma^{0}\right)^{2}=\mathbb{1}\,,$ $\left(\gamma^{i}\right)^{\dagger}=-\gamma^{i}$ $\implies\left(\gamma^{i}\right)^{2}=-\mathbb{1}\,,$ (185)
which imply that $\gamma^{0}$ is Hermitian and unitary, while $\gamma^{i}$ are anti-Hermitian and unitary.
Moreover, adopting the GR signature, the definition of the $\gamma^{5}$ matrix becomes
$\gamma^{5}=i\gamma^{0}\gamma^{1}\gamma^{2}\gamma^{3}=\begin{pmatrix}-\mathbb{1}&0\\ 0&\mathbb{1}\end{pmatrix}\,.$ (186)
Unlike other approaches found in the literature Bojowald:2010qpa ; Das:2008r , we do not modify the gamma matrices or the fermionic fields when transitioning to the curved spacetime formulation of the Dirac action. The key differences between the flat and curved spacetime cases arise from the introduction of the tetrad fields and a sign change in the anticommutation relations of the gamma matrices, as shown in Eq. (184). As a consequence, several familiar identities involving gamma matrices are altered. Below, we present a list of these modified identities, which can be verified straightforwardly
$\gamma^{I}\gamma_{I}=$ $-4\mathbb{1}\,,$ (187)
$\gamma^{I}\gamma^{J}\gamma_{I}=$ $2\gamma^{I}\,,$ (188)
$\gamma^{I}\gamma^{J}\gamma^{K}\gamma_{I}=$ $4\eta^{JK}\mathbb{1}\,,$ (189)
$\gamma^{I}\gamma^{J}\gamma^{K}\gamma^{L}\gamma_{I}=$ $2\gamma^{L}\gamma^{K}\gamma^{J}\,,$ (190)
$\gamma^{I}\gamma^{J}\gamma^{K}=$ $-\eta^{IJ}\gamma^{K}-\eta^{JK}\gamma^{I}+\eta^{IK}\gamma^{J}+i\epsilon^{LIJK}\gamma_{L}\gamma^{5}\,,$ (191)
$\epsilon^{IJKL}\sigma_{IJ}=$ $-2i\gamma^{5}\sigma^{KL}\,,$ (192)
$\gamma^{I}\sigma_{JK}+\sigma_{JK}\gamma^{I}=$ $\,i\epsilon^{I}\,_{JKL}\gamma_{5}\gamma^{L}\,,$ (193)
$\sigma_{JK}\gamma^{I}-\gamma^{I}\sigma_{JK}$ $=\eta_{J}^{I}\gamma_{K}-\eta_{K}^{I}\gamma_{J}\,.$ (194)
A.4 Dirac fermions in flat and curved spacetime
A general fermion field transforms according to a reducible representation [^5] A representation is called reducible if it can be decomposed into a direct sum of irreducible representations, each of which transforms independently under the group. $(\frac{1}{2},0)\oplus(0,\frac{1}{2})$. This means that it can be described in terms of two irreducible Weyl components: one left-chiral and one right-chiral. This decomposition underscores the fundamental role of Weyl fields, which serve as the basic building blocks from which any fermionic field can be constructed.
We begin by considering the case of flat spacetime. To construct a massive fermionic field, we consider the Lorentz-invariant combinations $\psi_{L}^{\dagger}\psi_{R}$ and $\psi_{R}^{\dagger}\psi_{L}$, which can be used to introduce a mass term in the Dirac Lagrangian:
$\mathcal{L}_{\text{mass}}=-m(\psi_{L}^{\dagger}\psi_{R}+\psi_{R}^{\dagger}\psi_{L})\,.$ (195)
When combined with the kinetic term, the resulting theory can describe Majorana fermions, provided the fields are real solutions of the Dirac equation. This would require imaginary gamma matrices. However, this scenario will not be explored further in this work. Instead, we will focus on generic complex solutions of the Dirac equation, without imposing any reality conditions. This leads us to the Dirac field, whose Lagrangian in terms of the Weyl components is given by
$\mathcal{L}_{\text{Dirac}}=i\psi_{R}^{\dagger}\sigma^{\mu}\partial_{\mu}\psi_{R}+i\psi_{L}^{\dagger}\tilde{\sigma}^{\mu}\partial_{\mu}\psi_{L}-m(\psi_{L}^{\dagger}\psi_{R}+\psi_{R}^{\dagger}\psi_{L})\,.$ (196)
The Dirac field is particularly useful for studying QED, which does not distinguish between left- and right-handed components, i.e., it is not a chiral theory. Since we are not concerned with analyzing chirality in our discussion, working in the standard (Dirac) representation is more convenient.
One of the future goals of this study will be to examine propagators in a modified version of QED. Importantly, the choice of representation for the gamma matrices (whether Dirac, Weyl, or Majorana) does not affect the physical structure or form of the propagators. Therefore, when referring to the Dirac field, we will adopt the four-component spinor formalism as defined in (176), and in this case, the Dirac Lagrangian takes the form
$\mathcal{L}_{\text{Dirac}}=i\overline{\Psi}\gamma^{\mu}\partial_{\mu}\Psi-m\overline{\Psi}\Psi\,,$ (197)
where the gamma matrices are given in Eq.(178).
To extend the previously discussed Lagrangian to curved spacetime, we follow the procedure outlined in Appendix A.3. The fermionic covariant derivative is introduced as defined in Eqs. (181)-(182). It is important to note that the mass term remains unchanged in curved spacetime, as it does not involve any explicit gamma matrix or derivative operators. As a result, the action for Dirac fields in a curved background takes the form
$S_{F}=\int d^{4}x|e|\left\{\frac{i}{2}\left(e_{I}^{a}\overline{\Psi}\gamma^{I}\mathfrak{D}_{a}\Psi-e_{I}^{a}\overline{\mathfrak{D}_{a}\Psi}\gamma^{I}\Psi\right)-m\overline{\Psi}\Psi\right\}\,.$ (198)
where all symbols carry their usual meaning.

### v2

@@ 5 Discussion and conclusions
In this work we have provided the full analysis of the canonical formulation of the system of gravity, fermions, and photons, and their interaction in a background-independent way. In particular, we have considered the interaction term of the CED, which to our knowledge, has not been considered in previous studies.
We have analyzed the constraint structure of the system in the Ashtekar-Barbero variables with fermionic field, and have shown how the connection variables are modified due to the presence of fermions. Furthermore, the expression for the Hamiltonian, spatial diffeomorphism and Guass constraints are given in a way that the contribution of each of the sectors, i.e., gravity, fermions, and photons, are clear. We have also derived the EOM of the matter sector, which reduce to the usual EOM of field theory on fixed curved or flat spacetime in the appropriate limits. Finally, the behavior of the system under parity transformation is studied in detail.
This analysis and the results pave the way for further canonical analysis of realistic matter in the universe involved in electromagnetic interactions in strong and dynamical (non-fixed) gravity regimes. Most importantly, this work opens the door for further detailed studies of an interactive system (CED in a dynamical background) using loop quantum gravity, polymer quantization, generalized uncertainty principle, or other theory that uses canonical quantization techniques. A rather immediate application of this analysis is in symmetry-reduced systems such as cosmological models involving electromagnetic matter, or collapse scenarios in spherical symmetric or rotating black holes, which up to now, have been done almost exclusively only with scalar fields.
@@ Acknowledgements.
Appendix A Brief complementary information on fermions
To remain consistent with standard QFT conventions Schwartz (2013); Weinberg (1995); Peskin and Schroeder (1995), we adopt the signature (+,,,)(+,-,-,-) in this appendix only. In the main text, we revert to the GR signature (,+,+,+)(-,+,+,+).
In curved spacetime, fermions couple to gravity through the spin connection $\omega_{\mu}^{IJ}$. The spinor covariant derivatives acting on $\Psi$ and its adjoint are defined as
$\mathfrak{D}_{\mu}\Psi$ $=\partial_{\mu}\Psi+\tfrac{1}{2}\omega_{\mu}^{IJ}\sigma_{IJ}\Psi,$ (137)
$\overline{\mathfrak{D}_{\mu}\Psi}$ $=\partial_{\mu}\overline{\Psi}+\tfrac{1}{2}\omega_{\mu}^{IJ}\overline{\Psi}\,\sigma_{IJ},$ (138)
@@ Appendix A Brief complementary information on fermions
$\sigma_{IJ}=\tfrac{1}{4}[\gamma_{I},\gamma_{J}],\,\{\gamma^{I},\gamma^{J}\}=-2\eta^{IJ},\,\gamma^{5}=i\gamma^{0}\gamma^{1}\gamma^{2}\gamma^{3}=\begin{pmatrix}-\mathbb{1}&0\\ 0&\mathbb{1}\end{pmatrix}$. (139)
No redefinition of the fermionic fields or gamma matrices is required when passing to curved spacetime; all gravitational effects enter through the tetrad and spin connection.
Appendix B First class nature of constraints
A constraint is said to be first class if its Poisson bracket with all other constraints vanishes weakly, i.e. on the constraint surface. Essentially if the Poisson bracket leads to a linear combination of other first class constraints. In order to show that the constraints given in Eqs. (45)-(47) are first class, we explicitly compute their mutual Poisson brackets.
For the constraints in this paper, one can show this by first evaluating the functional derivatives of each constraint with respect to the canonical variables $A_{c}^{p}$ and $E_{p}^{c}$,
$\frac{\delta\mathcal{G}_{i}}{\delta A_{c}^{p}}$ $=0\,$, (140)
$\frac{\delta\mathcal{G}_{i}}{\delta E^{c}_{p}}$ $=\frac{1}{\kappa}\varepsilon_{ij}\,^{k}K_{c}^{j}\,$, (141)
$\frac{\delta\mathcal{H}}{\delta A_{c}^{p}}$ $=\frac{1}{\kappa}\left[\partial_{a}\left(\frac{E^{c}_{i}E^{a}_{j}}{\sqrt{\text{det}h}}\varepsilon^{ij}\,_{p}\right)-\frac{1}{\sqrt{\text{det}h}}\varepsilon^{mn}\,_{p}(-E^{c}_{m}\varepsilon_{nk}\,^{l}A_{a}^{k}E^{a}_{l}+E^{a}_{m}\varepsilon_{nk}\,^{l}A_{a}^{k}E^{c}_{l})\right]$
$=\frac{1}{\sqrt{\text{det}h}}\varepsilon^{mn}\,_{p}(E^{c}_{m}\mathcal{D}_{a}E^{a}_{n}+E^{a}_{n}\mathcal{D}_{a}E^{c}_{m})\approx 0\,$, (142)
$\frac{\delta\mathcal{H}}{\delta E^{c}_{p}}$ $=\frac{1}{\kappa}\frac{1}{\sqrt{\text{det}\,h}}\epsilon^{ij}{}_{k}E^{a}_{i}\left(\mathcal{F}_{ac}^{k}(A)-(1+\beta^{2})\epsilon^{k}{}_{mn}K_{a}^{m}K_{c}^{n}-2\frac{1+\beta^{2}}{\beta}\mathcal{D}_{[a}^{(\Gamma)}K_{c]}^{k}\right)$
$-\frac{i}{2}\left(\overline{\Psi}\gamma^{p}\mathfrak{D}_{c}\Psi-\overline{\mathfrak{D}_{c}\Psi}\gamma^{p}\Psi\right)\,$, (143)
$\frac{\delta\mathcal{H}_{a}}{\delta A_{c}^{p}}$ $=\frac{1}{\beta\kappa}(-\partial_{a}E^{c}_{p}+\delta_{a}^{c}\partial_{b}E_{p}^{b}+\varepsilon^{j}\,_{pl}E^{b}_{j}A_{b}^{l}\delta_{a}^{c}+\varepsilon^{j}\,_{kp}E_{j}^{c}A_{a}^{k})\,$, (144)
$\frac{\delta\mathcal{H}_{a}}{\delta E^{c}_{p}}$ $=\frac{1}{\beta\kappa}(\mathcal{F}_{ac}^{p}(A)-(1+\beta^{2})\varepsilon^{p}\,_{kl}K_{a}^{k}K_{c}^{l})\,.$ (145)
In deriving these expressions, we have made repeated use of the compatibility condition $\mathcal{D}_{a}E^{a}_{i}\approx 0$. Using the above and the canonical Poisson brackets, the algebra of constraints reduce to
$\{\mathcal{G}_{i},\mathcal{H}\}$ $\approx 0\,$, (146)
$\{\mathcal{G}_{i},\mathcal{H}_{a}\}$ $=-\frac{1}{\beta\kappa^{2}}(-\varepsilon_{ij}\,^{p}K_{c}^{j}\mathcal{D}_{a}E^{c}_{p}+\varepsilon_{ij}\,^{p}K_{a}^{j}\mathcal{D}_{b}E^{b}_{p})\approx 0\,$, (147)
$\{\mathcal{H},\mathcal{H}_{a}\}$ $=\frac{1}{\beta\kappa^{2}}\frac{1}{\sqrt{\text{det}h}}\varepsilon^{ip}\,_{k}\left(\mathcal{F}_{ac}^{k}(A)-(1+\beta^{2})\epsilon^{k}{}_{mn}K_{a}^{m}K_{c}^{n}-2\frac{1+\beta^{2}}{\beta}\mathcal{D}_{[a}^{(\Gamma)}K_{c]}^{k}\right)$
$\times\left(-E^{a}_{i}\mathcal{D}_{a}E^{c}_{p}+E^{c}_{i}\mathcal{D}_{b}E_{p}^{b}\right)+\frac{i}{2\beta\kappa}\bigl[-\mathcal{D}_{a}E^{c}_{p}\left(\overline{\Psi}\gamma^{p}\mathfrak{D}_{c}\Psi-\overline{\mathfrak{D}_{c}\Psi}\gamma^{p}\Psi\right)$
$+\mathcal{D}_{b}E^{b}_{p}\left(\overline{\Psi}\gamma^{p}\mathfrak{D}_{a}\Psi-\overline{\mathfrak{D}_{a}\Psi}\gamma^{p}\Psi\right)\bigr]\approx 0\,$, (148)
which implies that that the Gauss, Hamiltonian, and diffeomorphism constraints close under the Poisson bracket and are all first class.